Problem 14-1 (Algorithmic)

Suppose that the R&B Beverage Company has a soft drink product that shows a constant annual demand rate of 3700 cases. A case of the soft drink costs R&B $2. Ordering costs are $18 per order and holding costs are 22% of the value of the inventory. R&B has 250 working days per year, and the lead time is 5 days. Identify the following aspects of the inventory policy:

  1. Economic order quantity. If required, round your answer to two decimal places.

    Q*=
  2. Reorder point.

    r =
  3. Cycle time. If required, round your answer to two decimal places.

    T = days
  4. Total annual cost. If required, round your answer to two decimal places.

    TC = $  
Answer #1

A.

Economic order quantity = (2*annual demand*ordering cost/holding cost per unit per year)^.5

Economic order quantity = (2*3700*18/(22%*2))^.5

Economic order quantity = 550.21 units or 550 units

B.

Reorder point = (3700/250)*5

Reorder point = 74

C.

Cycle time = 3700/550.21 = 6.72 or 7 days

D.

Total annual cost = (3700/550.21)*18 + (550.21/2)*(22%*2)

Total annual cost = $242.09

If cost of cases are also included, then

Total cost = 3700*2 + 242.09 = $7642.09